Gramps purchased a joint survivor annuity that pays $1,350 monthly over his remaining life and that of his wife, Gram. Gramps is 70 years old and Gram is 65 years old. Gramps paid $299,376 for the contract. How much income will Gramps recognize on the first payment?
Answer #1

this joint survivor annuity has an annual return multiple of 23.1 from table (vi) in Reg. sec.1.72-9.

the expected return is 1350*12*23.1 = $374,220 (annual payment * annual return multiple)

gramps paid $299,376 fr this contract

thus return on capital is 299376/374220 * 100 = 80%

hence 20% will be income

therefore first payment consists of 1350*20% = $270 of income